A bomb is dropped from a plane which is flying with a constant horizontal velocity. Locate the bomb after 1 second in relation to the plane.

1
Expert's answer
2021-12-30T11:59:19-0500

It’s directly below the plane.

Assuming that no speed has been lost by air drag.

Drag loss=0

Time(t)=1sec

it will continue at the same forward speed as the plane, but dropping lower with time.


We know that

Newton motion second law

"S=ut+\frac{1}{2}at^2\rightarrow(1)"

"u=0m/sec\\t=1sec\\g=9.8m/sec^2"

Equation (1) put value

"y=0+\frac{1}{2}\times9.8\times1^2\\y=\frac{1}{2}\times9.8\times1^2m\\y=4.9m\\"